Pratap Puri rowed 18 miles down the Delaware River in 2 hours, but the return trip took him 4 1/2 hours. Find the rate Pratap can row in still water, and find the rate of the current.
Let x= rate Pratap can row in still water and y = rate of the current
Downstream x+y
Upstream x-y
I think the first one is 2x+y=18
I can't figure out the second part. Something like:
x=y-4.5.

Answer by mananth(16949)

boat speed =x mph
current speed =y mph
against current x-y 4.50 hours
With current x+y 2.00 hours

Distance = same= 18 miles
t=d/r
18 / ( x - y )= 4.50
4.5 ( x - y ) = 18.00
4.5 x -4.5 y = 18 ....................1
18 / ( x + y )= 2.00
2.00 ( x + y ) = 18
2.00 x + 2.00 y = 18 ...............2
Multiply (1) by 2.00
Multiply (2) by 4.50
we get
9 x + -9 y = 36
9 x + 9 y = 81
18 x = 117
/ 18
x = 6.5 mph

plug value of x in (1)
4.5 x -4.5 y = 18
29.25 -4.5 y = 18
-4.5 y = 18 -29.25
-4.5 y = -11.25
y= 2.5 mph
boat 6.5 mph
current 2.5 mph
m.ananth@hotmail.ca